OpenStudy (anonymous):

The population of a village is 5000. It is increasing at the rate of 10% each year. At the end of 3 years, 1155 people migrate from the village. The net increase in population at the end of 3 years is? A) 27% B) 30% C) 9% D) 10%

OpenStudy (kira_yamato):

The population of a village is 5000. It is increasing at the rate of 10% each year. At the end of 3 years, 1155 people migrate from the village. The net increase in population at the end of 3 years is? The first part is GP with first term 5000 and common ratio 1.1. At the end of 3 years, before migration, population becomes 6050 (see attached for details) The next part you need to settle is migration. Since 1155 people migrate away, then no. of people after 3 years = 6050-1155 = 4895

OpenStudy (kira_yamato):

Wait... Is it immigrate into village or emigrate out of village

OpenStudy (kira_yamato):

If it is immigrate into village then different story. Same thing holds for part (i), but No. of people = 6050 + 1155 = 7205 Net growth = (7205-5000)/5000 *100% = 44.1% but there's no answer close....